OpenStudy (anonymous):

simplify 2sinx^7x/cos^3x *[cosx/2sinx]^3 please help

OpenStudy (anonymous):

\[(\frac{ 2\sin^7x }{ \cos^3x }) (\frac{ cosx }{ 2sinx })^3 \] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first simplify the cube \[(\frac{ 2\sin^7x }{ \cos^3x })(\frac{ \cos^3x }{ 2\sin^3x }) \] and do you know how to multiply fractions? i give you the general formula for it \[(\frac{ a }{ b })(\frac{ c }{ d }) = \frac{ ac }{ bd } \] multiply that for me

OpenStudy (anonymous):

2sin^7x*cos^3x/cos^3x*2sin^3x

OpenStudy (anonymous):

ok good now notice that \[\frac{\ 2cos^3x }{\ 2cos^3x } = 1 \] so you can cancel that out and you are left with \[\frac{ \sin^7x }{ \sin^3x }\]

OpenStudy (anonymous):

now notice that \[\sin^7x = (\sin^4x)(\sin^3x)\] so you can cancel out the denominator so whatever you are left with is the simplified form of the original equation

OpenStudy (anonymous):

so this is the final answer = 1+ sin^4xsin^3x

OpenStudy (anonymous):

uhhh its just \[\sin^4x \]

OpenStudy (anonymous):

oky i got it